# OmniAudio: Generating Spatial Audio from 360-Degree Video

### v1

@@ Appendix C Dataset Comparisons and Benchmark Construction
Several datasets have been developed to support research in 360-degree video and spatial audio processing. REC-STREET (Morgado et al., 2018), YT-ALL (Morgado et al., 2018), and YT-360 (Morgado et al., 2020) provide large-scale 360-degree video collections with First-order Ambisonics (FOA) audio (with total durations of 3.5, 113, and 246 hours, respectively), although their primary focus is on audio-visual correspondence learning. STARRS23 (Shimada et al., 2023), while supporting spatial audio generation from 360-degree videos, contains only 200 clips (7.5 hours), limiting its applicability for deep learning approaches. More recently, YT-Ambigen (Anonymous, 2024) introduced a dataset of 102K clips (totaling 142 hours), specifically designed for spatial audio generation; however, it is limited to fixed field-of-view (FoV) videos rather than full 360-degree content. Our Sphere360 dataset offers over 103K high-quality 360-degree video clips with FOA audio, totaling 288 hours in duration, specifically curated for spatial audio generation through meticulous data collection and cleaning pipelines. A comparison of Sphere360 with existing datasets is shown in Table 8.

### v3

@@ 6.2 Negative Impacts
6.3 Key Considerations
Regulation and Transparency: Developing guidelines and regulations to address the potential misuse of video-to-audio generation technology is essential.
Ethical Development: Researchers and developers should prioritize responsible development practices to mitigate potential harms and ensure
Acknowledgements
The research was supported by NSFC (No.62206234) from Mainland China, Early Career Scheme (ECS-HKUST22201322) from Hong Kong RGC, and National Natural Science Foundation of China under Grant No.62222211 and No.U24A20326.
@@ Appendix C Dataset Comparisons and Benchmark Construction
Several datasets (Morgado et al., 2020; Liu et al., 2023b) have been developed to support research in 360-degree video and spatial audio processing. REC-STREET (Morgado et al., 2018), YT-ALL (Morgado et al., 2018), and YT-360 (Morgado et al., 2020) provide large-scale 360-degree video collections with First-order Ambisonics (FOA) audio (with total durations of 3.5, 113, and 246 hours, respectively), although their primary focus is on audio-visual correspondence learning. STARRS23 (Shimada et al., 2023), while supporting spatial audio generation from 360-degree videos, contains only 200 clips (7.5 hours), limiting its applicability for deep learning approaches. More recently, YT-Ambigen (Anonymous, 2024) introduced a dataset of 102K clips (totaling 142 hours), specifically designed for spatial audio generation; however, it is limited to fixed field-of-view (FoV) videos rather than full 360-degree content. Our Sphere360 dataset offers over 103K high-quality 360-degree video clips with FOA audio, totaling 288 hours in duration, specifically curated for spatial audio generation through meticulous data collection and cleaning pipelines. A comparison of Sphere360 with existing datasets is shown in Table 8.
@@ Appendix G Additional Quantitative Results
More Analysis of Dual-Branch Strategies
To further investigate the effectiveness of our dual-branch approach, we conducted additional experiments comparing various FOV-cut strategies. Our original design employs a 120° frontal view to focus on prominent sound sources while incorporating panoramic features as contextual conditioning. We evaluated three alternative FOV-cut strategies to replace the local FOV video:
  • Hexadirectional 360° cuts (ERP+6cuts): Capturing views from front, back, left, right, up, and down directions.
  • Quadrant cuts (ERP+4cuts): Capturing views from front, left, right, and back directions.
  • Bipolar cuts (ERP+2cuts): Capturing views from front and back directions.
Table 11 presents the results of this comparison.
Table 11: Comparison of Dual-Branch Methods
Dual-Branch Method FD $\downarrow$ KL $\downarrow$ $\Delta_{abs}\theta$downarrow $\Delta_{abs}\phi$downarrow $\Delta_{Angular}$downarrow
ERP+Front 88.30 1.58 1.36 0.52 1.28
ERP+2FOV Cuts 95.84 1.77 1.37 0.55 1.30
ERP+4FOV Cuts 92.89 1.65 1.36 0.51 1.27
ERP+6FOV Cuts 90.16 1.59 1.37 0.52 1.26
The results demonstrate that our original dual-branch strategy (ERP+Front) still achieves the best audio quality, as reflected in the FD and KL metrics. Moreover, it achieves similar spatial metrics compared to the multiple FOV view approaches. This analysis supports our choice of using a single frontal view combined with equirectangular projection in our final model design.